5.Dubas Co. is a U.S.-based MNC that has a subsidiary in Germany and another subsidiary in Greece. Both subsidiaries frequently remit their earnings back to the parent company. The German subsidiary generated a net outflow of E2,000,000 this year, while the Greek subsidiary generated a net inflow of E1,500,000.What is the net inflow or outflow as measured in U.S. dollars this year? The exchange rate for the euro is $1.05
Answer #1

Net outflow is higher than net inflow hence there will be a Net outflow = Outflow-Inflow = 2000000-1500000 = E 500,000

Hence in USD it will be 500000*1.05 = $525000

Hence Net Outflow is USD = $525000